# Emotive Speech-to-Text Interfaces in XR: A Narrative Review of Psychophysiological and Accessibility Advances

### v2

@@ 4.3 Next Steps
    Can the integration of transcribed text in an Augmented Reality with corresponding visual elements such as pictures, emojis, and avatars enhance the conveyance of emotional expression in speech-to-text conversion?
    Strategy: Subsequently, conduct a user study to assess the effectiveness of this combined approach.
@@
    What is the optimal channel and format for delivering speech-to-text transcriptions while preserving the speaker’s facial expressions?
    Strategy: Investigate different text formatting options, including presenting a few words per line, highlighting emphasized words, and summarizing content while maintaining eye contact, to determine the most effective approach.
@@
    How can the user experience in speech-to-text technology be enhanced to minimize cognitive load for both the speaker and the user?
    Strategy: Explore strategies and design elements that contribute to an improved user experience in speech-to-text interactions.
@@
    Can the substitution of emotionally recognizable words with equivalent emojis and avatars, introduced dynamically in a live transcription, effectively preserve and convey emotional expression?
    Strategy: Examine the impact of replacing words with visual elements on emotional expression in speech-to-text, particularly in a live transcription setting.
@@ 5 Conclusion
In conclusion, this narrative literature review explored the evolving landscape of emotional expression in STT interfaces, with a focus on accessibility, psychophysiology, and immersive XR environments. A key insight from this review is that while emotion visualization in XR is gaining momentum—through ambient cues, avatars, and color metaphors—these techniques have not yet been coupled with real-time ASR captioning. This represents a crucial research gap, particularly for applications in education, healthcare, and collaborative virtual spaces.
This narrative review explores communication enhancement technologies for DHH individuals, including anyone who relies on language transcription services to help with communication, with a focus on the psychophysiology of STT interfaces in VR. The review highlights advancements in STT and closed captioning, emphasizing their relevance in enhancing communication accessibility, emphasizing how emotional expression is lost in the process of transcribing speech into text.
The results from selected studies matter as they not only contribute to the ongoing efforts to make communication technologies more inclusive for the DHH community but also pave the way for emotionally rich interactions in various contexts, from virtual environments to healthcare and education. The cited studies collectively suggest a growing recognition of the importance of emotional expression in human-computer interactions and the ongoing pursuit of innovative solutions to address this crucial aspect.
The integration of empathy into machines has the potential to redefine how we interact with technology, making these interactions more intuitive, adaptive, and emotionally resonant. However, ethical considerations, user preferences, and ongoing research are crucial in shaping the responsible and effective deployment of such technologies.
The review culminates in specific research recommendations that prioritize emotionally expressive STT interfaces, aiming to bridge accessibility with affective communication in immersive settings.
In conclusion, this synthesis provides a comprehensive overview of the evolution of communication enhancement technologies, spanning from live transcription and closed captioning to innovations in AR, VR, and speech emotion recognition. The interdisciplinary nature of the field is underscored, emphasizing the need for continued research and development to create more inclusive and accessible communication technologies for individuals with hearing impairments and the general language transcription user. As we navigate this dynamic landscape, the synthesis serves as a guide for future research directions and the development of impactful technologies that cater to diverse sensory and cognitive needs.

### v3

@@ 4.3 Future Work Directions
    Integrating Visual Augmentation in AR/VR STT Systems: Investigate how the use of emojis, animated avatars, and symbolic visuals can enhance emotional expression in real-time captions.
@@
    Designing Emotionally Resonant Caption Formats: Identify optimal caption formats and placement strategies that preserve speaker affect and reduce viewer distraction in immersive settings.
@@
    Reducing Cognitive Load through Personalization: Develop adaptive captioning interfaces that adjust visual emphasis, pacing, and layout based on user preferences and cognitive workload.
@@
    Dynamic Substitution of Emotive Language: Explore mechanisms for substituting emotionally expressive language with visual or symbolic equivalents (e.g., icons, colors) during live transcription.
  5. 5.
    Cross-Platform Evaluation Frameworks: Build comprehensive evaluation frameworks that assess the emotional clarity, usability, and accessibility of captioning strategies across XR platforms.
@@ 5 Conclusion
This narrative review explored the integration of emotional expression into STT interfaces within immersive XR environments, emphasizing the intersection of psychophysiology, accessibility, and user-centered design. While STT technologies have significantly advanced accessibility for DHH users, the omission of emotional sues in transcribed speech remains a critical barrier to natural and expressive communication.
Our findings highlight a growing recognition in HCI of the need for emotion-aware captioning, with emerging strategies such as emotion-driven avatars, color-coded overlays, and ambient cues offering promising—but largely disconnected—approaches. A key research gap lies in the lack of direct integration between real-time STT outputs and affective visualization techniques.
Addressing this gap requires interdisciplinary collaboration across affective computing, neuroscience, and XR systems design. Future research should investigate the effectiveness of real-time emotional augmentation—such as emojis, avatars, and dynamic text formatting—in reducing cognitive load and enhancing comprehension in AR/VR settings. This review highlights the urgent need to move beyond neutral captioning toward systems that can capture and convey the emotional context of speech.
By bridging accessibility with emotional expressiveness, the next generation of STT systems can support more inclusive, empathetic, and engaging communication—paving the way for richer user experiences in immersive environments and beyond.